Sales-Value-at-Split-off Method

Alomar Company manufactures four products from a joint production process: barlon, selene, plicene, and corsol. The joint costs for one batch are as follows:

Direct materials $69,000
Direct labor 35,000
Overhead 26,000

At the split-off point, a batch yields 1,000 barlon, 2,600 selene, 2,500 plicene, and 3,600 corsol. All products are sold at the split-off point: barlon sells for $17 per unit, selene sells for $23 per unit, plicene sells for $27 per unit, and corsol sells for $37 per unit.

Required:

Allocate the joint costs using the sales-value-at-split-off method. If required, round allocation rates to four decimal places and round the final allocations to the nearest dollar.

Allocated Joint Cost
Barlon $
Selene
Plicene
Corsol
Total $

(Note: The total of the allocated cost may not equal actual total costs to due to rounding.)
Answer #1

Allocation of Joint cost

Sales value at split off Allocated Joint cost
Barlon 1000*17 = 17000 0.0613 7969
Selene 2600*23 = 59800 0.2155 28015
Plicene 2500*27 = 67500 0.2432 31616
Corsol 3600*37 = 133200 0.48 62400
Total 277500 130000